OpenStudy (anonymous):

find a polynomial f(x) of degree 3 with real coefficients and the following zeros -1,2i

OpenStudy (anonymous):

anyone?

Directrix (directrix):

Use the Factor Theorem: Factor Theorem, which states: If f(a) = 0, then (x-a) is a factor of f(x). So, if -1 is a zero, then (x - (-1)) = (x + 1) is a factor. And, if 2i is a zero, then its conjugate, -2i, is a zero. ^^ Apply the Factor Theorem to them, also.l

OpenStudy (anonymous):

the answer here isn't very clear

Directrix (directrix):

It is not an answer. It is a direction for your thinking. In short, f(x) = (x + 1) * (x + 2i ) * (x - 2i ).